# PACMANN: Point Adaptive Collocation Method for Artificial Neural Networks

### v1

@@ Abstract
Physics-Informed Neural Networks (PINNs) are an emerging tool for approximating the solution of Partial Differential Equations (PDEs) in both forward and inverse problems. PINNs minimize a loss function which includes the PDE residual determined for a set of collocation points. Previous work has shown that the number and distribution of these collocation points have a significant influence on the accuracy of the PINN solution. Therefore, the effective placement of these collocation points is an active area of research. Specifically, adaptive collocation point sampling methods have been proposed, which have been reported to scale poorly to higher dimensions. In this work, we address this issue and present the Point Adaptive Collocation Method for Artificial Neural Networks (PACMANN). Inspired by classic optimization problems, this approach incrementally moves collocation points toward regions of higher residuals using gradient-based optimization algorithms guided by the gradient of the squared residual. We apply PACMANN for forward and inverse problems, and demonstrate that this method matches the performance of state-of-the-art methods in terms of the accuracy/efficiency tradeoff for the low-dimensional problems, while outperforming available approaches for high-dimensional problems; the best performance is observed for the Adam optimizer. Key features of the method include its low computational cost and simplicity of integration in existing physics-informed neural network pipelines.
@@
Differential equations , Physics-informed neural networks (PINNs) , Adaptive sampling , Residual gradient
journal: arXiv
@@
[c]organization=Delft Institute for Computational Science and Engineering, addressline=Delft University of Technology, state=2628 CD, country=The Netherlands
@@ 1 Introduction
Physics-Informed Neural Networks (PINNs) build upon the ability of deep neural networks to serve as universal function approximators, as established by Hornik et al. [1] in 1989. Based on this finding, several methods were developed to solve Ordinary Differential Equations (ODEs) and Partial Differential Equations (PDEs) using neural networks, originally proposed by [2, 3]. Supported by these developments and recent advances in computational tools, notably automatic differentiation in 2015 [4], Raissi et al. [5] proposed the name and framework of physics-informed neural networks in 2019 and their use to approximate the solution of PDEs in both forward and inverse problems. In their classic form, PINNs approximate the solution by minimizing a loss function based on boundary conditions, initial conditions, and the PDE residual sampled over a set of collocation points. In the following years, PINNs have been applied in a variety of fields [6, 7], such as fluid dynamics [8, 9, 10], heat transfer [11, 12], material sciences [13, 14] and electromagnetism [15, 16].
@@
In 2020, Mao et al. [8] explored the impact of collocation point placement on prediction accuracy for solutions with discontinuities. They demonstrated that, when discontinuities are known a priori, manually increasing the density of points near these regions improves prediction accuracy. To address scenarios where solution features are unknown before training, adaptive algorithms for collocation point selection were developed. For instance, Lu et al. [17] introduced Residual-based Adaptive Refinement (RAR) in 2021, the first adaptive sampling algorithm, which places additional points in regions with the largest PDE residuals. RAR improves prediction accuracy, capturing features such as a discontinuity in a Burgers’ equation problem better than a static grid. Instead of sampling points only where the residual is the largest, Nabian et al. [18] proposed to randomly resample all points in the domain based on a Probability Density Function (PDF) proportional to the residual. This approach samples a higher density of points in high residual areas, resulting in accelerated convergence of the PINN. Based upon these studies, Wu et al. [19] presented two additional sampling algorithms, Residual-based Adaptive Distribution (RAD) and Residual-based Adaptive Refinement with Distribution (RAR-D). In RAD, all collocation points are resampled using a PDF based on the residual. RAR-D is a combination of RAR and RAD, where collocation points are sampled in addition to the existing collocation points according to the same probability density function used by RAD. Both approaches lead to a higher prediction accuracy, specifically for PDEs with complex solutions due to, for example, steep gradients. Moreover, RAD was found to outperform the method proposed by Nabian et al. [18].
@@
Several adaptations to the aforementioned PDF-based sampling algorithms have been proposed. Guo et al. [20], for example, state that adaptive sampling methods should obey temporal causality. They propose the adaptive causal sampling method, which decomposes the domain into subdomains where the ratio of points sampled in each subdomain is based on the PDE residual and a temporal weight, ensuring temporal causality. This approach was found to enhance the prediction accuracy and computational efficiency of PINNs in problems with nonlinear PDEs containing higher-order derivatives. Furthermore, Mao et al. [21] consider the gradient of the solution by sampling additional points in subdomains with large residuals and large solution gradients. Liu et al. [22] propose to add points with large residual gradients to the set of collocation points used for training. Both Mao et al. and Liu et al. report an improvement in accuracy for problems with solutions exhibiting steep gradients.
While the aforementioned collocation point sampling methods have proven effective in low-dimensional problems, these approaches to resampling are computationally expensive for high-dimensional problems as reported by Wu et al. [19]. Methods such as RAD or RAR are computationally expensive for these problems due to the cost of evaluating the residual in a sufficiently large number of points, either to construct the probability density function or to identify additional points for inclusion in the training process.
@@
Other approaches have been proposed to sample collocation points in high-dimensional problems. For instance, Tang et al. [23] propose the DAS-PINNs approach, which samples according to a Deep Adaptive Sampling (DAS) method and uses KRnet [24], a deep generative model, to approximate the PDF proportional to the residual. However, this approach is complex to implement in existing PINNs pipelines due to its dual-network framework.
@@
The present work aims to develop a collocation point resampling method that efficiently scales to higher dimensions without introducing significant computational overhead while maintaining the advantages of previous approaches. We propose the Point Adaptive Collocation Method for Artificial Neural Networks (PACMANN), which uses the gradient of the squared residual to move collocation points toward areas with a large residual. In this approach, collocation point resampling is formulated as a maximization problem of the squared residual, where established optimization methods may be employed to move points. PACMANN includes two main hyperparameters, the stepsize and the number of steps taken by the optimization algorithm.
@@
First, we investigate the performance of PACMANN in combination with a variety of optimization algorithms for two low-dimensional problems: a 1D Burgers’ equation and a 1D Allen-Cahn equation. We then perform sensitivity studies for the number of collocation points, the number of iterations between resampling, and the two aforementioned hyperparameters of our proposed method. In addition, we demonstrate the suitability of PACMANN to high-dimensional and inverse problems: a 5D Poisson’s equation and a 2D inverse Navier-Stokes equation. For all problems under consideration, we compare the performance of our method in terms of prediction accuracy and computational cost to state-of-the-art adaptive and non-adaptive sampling methods. Notably, our results show that our method matches the performance of state-of-the-art methods in terms of the accuracy/efficiency tradeoff for low-dimensional problems and efficiently scales to high-dimensional problems, where it outperforms state-of-the-art methods.
@@
This paper is organized as follows. In Section 2, a review of the PINNs framework is given, followed by a description of PACMANN and the optimization algorithms considered in this study. Next, in experimental studies in Section 3, we compare the accuracy and computational cost of PACMANN to other sampling methods for three forward problems and an inverse problem under varying hyperparameters. Finally, in Section 4, we summarize our findings.
@@ 2 Methodology
This section begins with a brief review of PINNs based on the framework presented by Raissi et al. [5] in 2019. Next, we propose the novel PACMANN, followed by a summary of the optimization algorithms considered in this work.
2.1 Physics-Informed Neural Networks
PINNs approximate the solution of PDEs using neural networks. Generally, we consider PDEs of the form
$\boldsymbol{u_{t}}+\mathcal{N}[\boldsymbol{u}]=0,\quad\boldsymbol{x}\in\boldsymbol{\Omega},\quad t\in[0,T],$
@@ 2.1 Physics-Informed Neural Networks
$\boldsymbol{u}\left(\boldsymbol{x},0\right)$ $=h\left(\boldsymbol{x}\right),\quad\boldsymbol{x}\in\boldsymbol{\Omega},$
$\mathcal{B}[\boldsymbol{u}]\left(\boldsymbol{x},t\right)$ $=0,\quad\boldsymbol{x}\in\partial\boldsymbol{\Omega},\quad t\in[0,T],$
where $\mathcal{N}[\cdot]$ is a linear or nonlinear differential operator, and $\mathcal{B}[\cdot]$ is a boundary operator corresponding to a set of boundary conditions. In addition, $\boldsymbol{x}\in\mathbb{R}^{d}$ and $t\in[0,T]$ denote the spatial and temporal coordinates, $\boldsymbol{\Omega}\subset\mathbb{R}^{d}$, $\partial\boldsymbol{\Omega}$ forms the boundary. The set of collocation points $\{\boldsymbol{x}_{r}^{i},t_{r}^{i}\}_{i=1}^{N_{r}}$ in the domain $\mathcal{D}$ are represented by $\boldsymbol{X}$. Furthermore, $\boldsymbol{u}$ denotes the exact solution.
The PINN consists of a (deep) neural network with the coordinates $(\boldsymbol{x},t)$ as inputs and $\hat{\boldsymbol{u}}(\boldsymbol{x},t,\boldsymbol{\theta})$ as output. The trainable parameters $\boldsymbol{\theta}$ of this neural network are trained by minimizing the loss function $\mathcal{L}(\boldsymbol{\theta})$
$\boldsymbol{\theta}^{*}=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathcal{L}\left(\boldsymbol{\theta}\right).$
@@
$\mathcal{L}_{r}\left(\boldsymbol{\theta}\right)=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left(\hat{\boldsymbol{u_{t}}}\left(\boldsymbol{x}_{r}^{i},t_{r}^{i},\boldsymbol{\theta}\right)+\mathcal{N}[\hat{\boldsymbol{u}}]\left(\boldsymbol{x}_{r}^{i},t_{r}^{i},\boldsymbol{\theta}\right)\right)^{2},$
$\mathcal{L}_{ic}\left(\boldsymbol{\theta}\right)=\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left(\hat{\boldsymbol{u}}\left(\boldsymbol{x}_{ic}^{i},0,\boldsymbol{\theta}\right)-h\left(\boldsymbol{x}_{ic}^{i}\right)\right)^{2},$
$\mathcal{L}_{bc}\left(\boldsymbol{\theta}\right)=\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left(\mathcal{B}[\hat{\boldsymbol{u}}]\left(\boldsymbol{x}_{bc}^{i},t_{bc}^{i},\boldsymbol{\theta}\right)\right)^{2}.$
Here, $\mathcal{L}_{r}$, $\mathcal{L}_{ic}$, and $\mathcal{L}_{bc}$ represent the loss terms for the PDE residual, initial conditions, and the boundary conditions, respectively. Furthermore, $N_{r}$, $N_{ic}$, and $N_{bc}$ denote the number of collocation points of the aforementioned terms. The hyperparameters $\lambda_{r}$, $\lambda_{ic}$, and $\lambda_{ic}$ are scalar weights used to balance the loss function. Each loss term is evaluated over a set of data points, where $\{\boldsymbol{x}_{r}^{i},t_{r}^{i}\}_{i=1}^{N_{r}}$ is a set of collocation points located in the interior of the domain, $\{\boldsymbol{x}_{ic}^{i}\}_{i=1}^{N_{ic}}$ is a set of points sampled at the initial time, and $\{\boldsymbol{x}_{bc}^{i},t_{bc}^{i}\}_{i=1}^{N_{bc}}$ is a set sampled along the boundary. These points may be fixed during training, resampled through periodic random resampling, or resampled using adaptive sampling methods based on guiding information, such as the PDE residual.
To train the model parameters $\boldsymbol{\theta}$, the gradient of the loss function with respect to the parameters is determined using back-propagation [25]. Next, the model parameters are updated with an optimization algorithm, often based on the gradient descent method, such as the Adam optimizer [26]. Similarly, the gradients of $\boldsymbol{u}\left(\boldsymbol{x},t,\boldsymbol{\theta}\right)$ with respect to the input coordinates $\left(\boldsymbol{x},t\right)$ as required by $\mathcal{N}[\cdot]$ and potentially $\mathcal{B}[\cdot]$ in equation (2.1) are computed using automatic differentiation.
For problems that incorporate reference data during training, such as inverse problems, an additional loss term $\mathcal{L}_{ref}$ is added to the loss function described by equation (1), where
@@
This term corresponds to the mean squared error between the (noisy) observed data $\boldsymbol{u}_{ref}$ at the set of data points $\{\boldsymbol{x}^{i},t^{i}\}_{i=1}^{N_{ref}}$ and the approximation $\hat{\boldsymbol{u}}\left(\boldsymbol{x}^{i},t^{i},\boldsymbol{\theta}\right)$ given by the neural network. The addition of this term leads to the following loss function
@@
where $\lambda_{ref}$ is the scalar weight assigned to the reference data loss term.
2.2 Point Adaptive Collocation Method for Artificial Neural Networks
In this work, we propose PACMANN which uses the gradient of the squared residual as guiding information in combination with established optimization algorithms to gradually move collocation points toward areas of high residuals; see Algorithm 1 and Figure 1. This method turns collocation point resampling into a maximization problem. Training the model parameters $\boldsymbol{\theta}$ can thus be formulated as the following min-max problem
$\boldsymbol{\theta}^{*}=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\left[\lambda_{ic}\mathcal{L}_{ic}\left(\boldsymbol{\theta}\right)+\lambda_{bc}\mathcal{L}_{bc}\left(\boldsymbol{\theta}\right)+\lambda_{ref}\mathcal{L}_{ref}\left(\boldsymbol{\theta}\right)+\max_{\boldsymbol{X}\subset\mathcal{D}}\lambda_{r}\mathcal{L}_{r}\left(\boldsymbol{X},\boldsymbol{\theta}\right)\right].$
Where only the collocation points are moved, while points such as those sampled along the boundary are fixed in place throughout training. This approach ensures that a sufficient number of points are placed along the initial and boundary conditions to accurately compute the respective loss terms.
First, $N_{r}$ collocation points are sampled using a uniform sampling method, such as an equispaced uniform grid or the Hammersley sequence [27]. The PINN is then trained on this set of collocation points for a number of iterations determined by the resampling period $P$. Next, training is stopped and the gradients of the squared residual $r^{2}(\boldsymbol{x},t)$ with respect to the input coordinates, given by $\nabla_{\boldsymbol{x}}r^{2}(\boldsymbol{x},t)$ and $\frac{\partial}{\partial t}r^{2}(\boldsymbol{x},t)$, are determined for each collocation point $\left(\boldsymbol{x},t\right)$. The collocation points are subsequently moved in the direction of increasing residual based on the direction and magnitude of the residual gradient, as described by
[…]
Here, $s$ determines the stepsize of each move. Since training is stopped, the residual landscape is static and the residual gradient can be determined again for the new location of the collocation point, allowing the process to be repeated. Therefore, collocation points can be moved several times, given by the number of steps $T$. If a point moves outside the domain, it is removed and a replacement point is sampled in the domain based on a uniform probability distribution. Next, the neural network is trained for another $P$ iterations, after which the process of moving collocation points is repeated.
1 Sample a set $\boldsymbol{X}$ of $N_{r}$ collocation points $\{\boldsymbol{x}_{r}^{i},t_{r}^{i}\}_{i=1}^{N_{r}}$ using a uniform sampling method;

### v2

@@ Abstract
Physics-Informed Neural Networks (PINNs) have emerged as a tool for approximating the solution of Partial Differential Equations (PDEs) in both forward and inverse problems. PINNs minimize a loss function which includes the PDE residual determined for a set of collocation points. Previous work has shown that the number and distribution of these collocation points have a significant influence on the accuracy of the PINN solution. Therefore, the effective placement of these collocation points is an active area of research. Specifically, available adaptive collocation point sampling methods have been reported to scale poorly in terms of computational cost when applied to high-dimensional problems. In this work, we address this issue and present the Point Adaptive Collocation Method for Artificial Neural Networks (PACMANN). PACMANN incrementally moves collocation points toward regions of higher residuals using gradient-based optimization algorithms guided by the gradient of the PINN loss function, that is, the squared PDE residual. We apply PACMANN for forward and inverse problems, and demonstrate that this method matches the performance of state-of-the-art methods in terms of the accuracy/efficiency tradeoff for the low-dimensional problems, while outperforming available approaches for high-dimensional problems. Key features of the method include its low computational cost and simplicity of integration into existing physics-informed neural network pipelines. The code is available at https://github.com/CoenVisser/PACMANN.
@@
Differential equations , Physics-informed neural networks (PINNs) , Adaptive sampling , Collocation points , Residual gradient
journal: Journal of Computational Physics
@@
[c]organization=Delft Institute for Computational Science and Engineering, addressline=Delft University of Technology, state=2628 CD Delft, country=The Netherlands
@@ 1 Introduction
Physics-Informed Neural Networks (PINNs) build upon the ability of deep neural networks to serve as universal function approximators, as established by Cybenko [1] and Hornik et al. [2] in 1989. Based on these findings, several methods were developed to solve Ordinary Differential Equations (ODEs) and Partial Differential Equations (PDEs) using neural networks, originally proposed by [3, 4]. Supported by these developments and recent advances in computational tools, notably automatic differentiation in 2015 [5], Raissi et al. [6] proposed the name and framework of Physics-Informed Neural Networks and their use to approximate the solution of PDEs in both forward and inverse problems; their work was published in 2019. Since then, PINNs have been applied in a variety of fields [7, 8], such as fluid dynamics [9, 10, 11], heat transfer [12, 13], material sciences [14, 15], and electromagnetism [16, 17].
@@
In their classic form, PINNs approximate the solution of differential equations by minimizing a loss function incorporating boundary conditions, initial conditions, and the PDE residual sampled over a set of collocation points. In 2020, Mao et al. [9] explored the impact of collocation point placement on prediction accuracy for solutions exhibiting discontinuities. They demonstrated that, when discontinuities are known a priori, manually increasing the density of points near these regions improves prediction accuracy. To address scenarios where solution features are unknown before training, adaptive algorithms for collocation point selection were developed. For instance, Lu et al. [18] introduced Residual-based Adaptive Refinement (RAR) in 2021, the first adaptive sampling algorithm, which places additional points in regions with the largest PDE residuals. RAR improves prediction accuracy, capturing features such as a discontinuity better than a static grid for the Burgers’ equation. Instead of sampling points only where the residual is the largest, Nabian et al. [19] proposed to randomly resample all points in the domain based on a Probability Density Function (PDF) proportional to the loss function. This approach samples a higher density of points in high residual areas, resulting in accelerated convergence of the PINN. Based upon these studies, Wu et al. [20] presented two additional sampling algorithms, Residual-based Adaptive Distribution (RAD) and Residual-based Adaptive Refinement with Distribution (RAR-D). In RAD, all collocation points are resampled using a PDF defined by a nonlinear function of the PDE residual. RAR-D is a combination of RAR and RAD, where collocation points are sampled in addition to the existing ones according to the same probability density function used by RAD. Both approaches lead to a higher prediction accuracy, specifically for PDEs with complex solutions due to, for example, steep gradients. Moreover, RAD was found to outperform the method proposed by Nabian et al. [19].
@@
Several adaptations to the aforementioned PDF-based sampling algorithms have been proposed. Guo et al. [21], for example, propose an adaptive causal sampling method, which decomposes the domain into subdomains where the ratio of points sampled in each subdomain is based on the PDE residual and a temporal weight, ensuring temporal causality. This approach was found to enhance the prediction accuracy and computational efficiency of PINNs in problems with nonlinear PDEs containing higher-order derivatives. Furthermore, Mao et al. [22] consider the gradient of the solution by sampling additional points in subdomains with large residuals and large solution gradients. Liu et al. [23] propose to add points with large residual gradients to the set of collocation points used for training. Both Mao et al. and Liu et al. report an improvement in accuracy for problems with solutions exhibiting steep gradients.
@@
While the aforementioned collocation point sampling methods have proven effective in low-dimensional problems, these approaches to resampling are computationally expensive for high-dimensional problems, as reported by Wu et al. [20]. Specifically, RAD or RAR are computationally expensive for these problems due to the cost of evaluating the residual at a sufficiently large number of points, either to construct the probability density function or to identify additional points for inclusion in the training process. Other approaches have been proposed to sample collocation points in high-dimensional problems. For instance, Tang et al. [24] propose the DAS-PINNs approach, which samples according to a Deep Adaptive Sampling (DAS) method and uses KRnet [25], a deep generative model, to approximate the PDF proportional to the residual. However, this approach is not straightforward to integrate into existing PINNs pipelines due to its dual-network framework.
@@
In this work, we present a collocation point resampling method that scales to higher dimensions more efficiently without introducing significant computational overhead while maintaining the accuracy improvements achieved by previous approaches. We propose the Point Adaptive Collocation Method for Artificial Neural Networks (PACMANN), which uses the gradient of the squared residual to move collocation points toward areas with a large residual. In this approach, collocation point resampling is formulated as a maximization problem of the squared residual. First, the PINN is trained on a static grid of collocation points. After a certain number of iterations, this process is paused and the gradient of the squared residual is determined for the input coordinates of each collocation point. Based on the magnitude and direction provided by these gradients, points are moved to maximize the squared residual using established optimization methods. Since the residual landscape is static while training is paused, the process of moving points may be repeated iteratively. PACMANN includes four main hyperparameters: the resampling period, the optimizer for moving the collocation points, the stepsize, and the number of steps taken by the optimization algorithm. Key features of the method include its low computational cost and simplicity of integration in existing physics-informed neural network pipelines. Our approach builds on the work of Wang et al. [26], who, independently of the aforementioned developments, found that iteratively updating the placement of collocation points by applying gradient ascent over the $L^{\infty}$ physics-informed loss results in a greater prediction accuracy for the Hamilton-Jacobi-Bellman equation.
@@
First, we investigate the performance of PACMANN in combination with a variety of optimization algorithms for two low-dimensional problems: the one-dimensional Burgers’ and Allen-Cahn equations. We then perform sensitivity studies on the number of collocation points and the method’s hyperparameters. In addition, we demonstrate the suitability of PACMANN for high-dimensional and inverse problems. As test cases, we consider an inverse problem based on the two-dimensional Navier-Stokes equation, the Poisson’s equation in five dimensions, and the three-dimensional Navier-Stokes equation. Finally, we apply PACMANN to a problem involving a re-entrant corner in a disk, showing the effectiveness of our approach on irregular domains. For all problems under consideration, we compare the performance of our method in terms of prediction accuracy and computational cost to state-of-the-art adaptive and non-adaptive sampling methods. Notably, our results show that our method matches the performance of state-of-the-art methods in terms of the accuracy/efficiency tradeoff for low-dimensional problems while efficiently scaling to high-dimensional problems, where it outperforms state-of-the-art methods.
@@
This paper is organized as follows: In Section 2, we briefly review the PINNs framework, followed by a description of PACMANN. Next, in experimental studies in Section 3, we compare the accuracy and computational cost of PACMANN to other state-of-the-art sampling methods for five forward problems and an inverse problem. Finally, in  Section 4, we summarize our findings.
@@ 2 Methodology
This section begins with a brief review of PINNs based on the framework presented by Raissi et al. [6] in 2019. Next, we propose the novel PACMANN.
2.1 Physics-Informed Neural Networks (PINNs)
PINNs approximate the solution of PDEs using neural networks. Generally, we consider PDEs of the form: find $\boldsymbol{u}$ such that
$\boldsymbol{u_{t}}+\mathcal{N}[\boldsymbol{u}]=0,\quad\boldsymbol{x}\in\boldsymbol{\Omega},\quad t\in[0,T],$ (1)
@@ 2.1 Physics-Informed Neural Networks (PINNs)
$\boldsymbol{u}\left(\boldsymbol{x},0\right)$ $=h\left(\boldsymbol{x}\right),$ $\boldsymbol{x}\in\boldsymbol{\Omega},$
$\mathcal{B}[\boldsymbol{u}]\left(\boldsymbol{x},t\right)$ $=0,$ $\boldsymbol{x}\in\partial\boldsymbol{\Omega},\quad t\in[0,T],$
where $\mathcal{N}[\cdot]$ is a linear or nonlinear differential operator, and $\mathcal{B}[\cdot]$ is a boundary operator corresponding to a set of boundary conditions. In addition, $\boldsymbol{x}\in\boldsymbol{\Omega}\subset\mathbb{R}^{d}$ and $t\in[0,T]$ denote the spatial and temporal coordinates, respectively, and we write $\partial\boldsymbol{\Omega}$ for the boundary of $\boldsymbol{\Omega}$. We denote the space-time domain by $\mathcal{D}\coloneqq\overline{\boldsymbol{\Omega}}\times[0,T]$.
The PINN consists of a (deep) neural network with the coordinates $(\boldsymbol{x},t)$ as inputs and $\hat{\boldsymbol{u}}(\boldsymbol{x},t,\boldsymbol{\theta})$ as output, approximating $\boldsymbol{u}(\boldsymbol{x},t)$. The trainable parameters $\boldsymbol{\theta}$ of this neural network are trained by minimizing a specific loss function $\mathcal{L}(\boldsymbol{\theta})$:
$\boldsymbol{\theta}^{*}=\operatorname*{arg\,min}_{\boldsymbol{\theta}}\mathcal{L}\left(\boldsymbol{\theta}\right).$ (2)
@@
$\mathcal{L}_{r}\left(\boldsymbol{\theta}\right)=\frac{1}{N_{r}}\sum_{i=1}^{N_{r}}\left(\hat{\boldsymbol{u_{t}}}\left(\boldsymbol{x}_{r}^{i},t_{r}^{i},\boldsymbol{\theta}\right)+\mathcal{N}[\hat{\boldsymbol{u}}]\left(\boldsymbol{x}_{r}^{i},t_{r}^{i},\boldsymbol{\theta}\right)\right)^{2},$ (4)
$\mathcal{L}_{ic}\left(\boldsymbol{\theta}\right)=\frac{1}{N_{ic}}\sum_{i=1}^{N_{ic}}\left(\hat{\boldsymbol{u}}\left(\boldsymbol{x}_{ic}^{i},0,\boldsymbol{\theta}\right)-h\left(\boldsymbol{x}_{ic}^{i}\right)\right)^{2},$ (5)
$\mathcal{L}_{bc}\left(\boldsymbol{\theta}\right)=\frac{1}{N_{bc}}\sum_{i=1}^{N_{bc}}\left(\mathcal{B}[\hat{\boldsymbol{u}}]\left(\boldsymbol{x}_{bc}^{i},t_{bc}^{i},\boldsymbol{\theta}\right)\right)^{2}$ (6)
represent the loss terms for the PDE residual, initial conditions, and the boundary conditions, respectively. Furthermore, $N_{r}$, $N_{ic}$, and $N_{bc}$ denote the numbers of collocation points of the aforementioned terms. The hyperparameters $\lambda_{r}$, $\lambda_{ic}$, and $\lambda_{bc}$ are scalar weights used to balance the loss function. Each loss term is evaluated over a set of data points, where $\{(\boldsymbol{x}_{r}^{i},t_{r}^{i})\}_{i=1}^{N_{r}}$ is a set of collocation points located in the interior of the domain, $\{(\boldsymbol{x}_{ic}^{i},0)\}_{i=1}^{N_{ic}}$ is a set of points sampled at the initial time, and $\{(\boldsymbol{x}_{bc}^{i},t_{bc}^{i})\}_{i=1}^{N_{bc}}$ is a set sampled along the boundary. These points may be fixed during training [20], resampled through periodic random resampling [20], or resampled using adaptive sampling methods based on guiding information, such as the PDE residual [18, 19, 20]. Note that we assume sufficient regularity, existence of a strong-form solution of Equation 1, for the PINN loss function to be meaningful.
To train the model parameters $\boldsymbol{\theta}$, the gradient of the loss function with respect to the parameters is determined using back-propagation [27]. Next, the model parameters are updated with an optimization algorithm, often based on the gradient descent method, such as the Adam optimizer [28]. Similarly, the derivatives of $\hat{\boldsymbol{u}}\left(\boldsymbol{x},t,\boldsymbol{\theta}\right)$ with respect to the input coordinates $\left(\boldsymbol{x},t\right)$ as required by $\mathcal{N}[\cdot]$ and potentially $\mathcal{B}[\cdot]$ in Equations 4 and 6 are computed using automatic differentiation; the initial conditions and the corresponding loss function Equation 5 may also depend on the temporal derivative of $\hat{\boldsymbol{u}}$, but we omit these cases for simplicity.
For problems that incorporate reference data during training, such as inverse problems, an additional loss term $\mathcal{L}_{ref}$ is added to the loss function described by Equation 3, where
@@
This term corresponds to the mean squared error between the (noisy) observed data $\boldsymbol{u}_{ref}$ at the set of data points $\{(\boldsymbol{x}^{i},t^{i})\}_{i=1}^{N_{ref}}$ and the approximation $\hat{\boldsymbol{u}}\left(\boldsymbol{x}^{i},t^{i},\boldsymbol{\theta}\right)$ given by the neural network. Adding this term leads to the following loss function
@@
where $\lambda_{ref}$ is the additional scalar weight assigned to the reference data loss term.
2.2 Point Adaptive Collocation Method for Artificial Neural Networks (PACMANN)
1Sample a set $\boldsymbol{X_{r}}$ of $N_{r}$ collocation points $\{\boldsymbol{x}_{r}^{i},t_{r}^{i}\}_{i=1}^{N_{r}}$ using a uniform sampling method;